Write an equation of a line with a slope of 0 and passing through the point (5,4).
Answer by jrfrunner(365)

any line with slope=0 is a flat line
which means it is in the form: y=b
---
if you look at the slope intercept form of a line equation
y=m*x+b , m=slope
If m=0 then y=0*x+b=b
--
So it will include all the points which have y=b
--
In your case, since it must include the point (5,4) and its y component is 4, The line with 0 slope containing this point must be y=4